Cryogenic Distribution System
All major leaks have been repaired and preparations for the upcoming cooldown are now under way.
The Cryogenic Distribution System (CDS) is a 310-meter-long piping system which delivers the cooling power to all cryomodules in the superconducting linac by means of a constant flow of supercritical helium at 4.5 Kelvin and 3 bar.
CDS commissioning is a prerequisite for the start of cryomodule installation.
Linac Warm Units
The first Linac Warm Unit (LWU) for the Spoke section has been transported to the tunnel.
The magnets in the LWUs are quadrupoles which will focus and compress the beam both vertically and horizontally. ESS will have 71 LWU's, placed in-between the cryomodules and in the last part of the tunnel, where the protons are bent towards the Target.
Drift Tube Linac (DTL)
Installation of DTL3 is ongoing in the Normal Conducting Linac. DTL4 is under vacuum tests.
Bead pull tests were completed for DTL2, which is now on the trolley made ready for the installation of water manifolds. The plan is to transport it to the tunnel in mid-November.
12 out of 26 tetrode-based RF power stations are in place in the Spoke cells, with power soak tests ongoing.
All klystrons are in place in the medium beta section of the Gallery, with four klystrons on each side of the cell. Two modulators will be placed in cell 190, to complete the medium beta RF cell set-up. Power tests continue.
The third high beta cryomodule (CM33) has been delivered to ESS from CEA in France.
High beta cryomodule CM31 is inside the Test Stand 2 bunker getting ready for cooldown, and CM32 is parked outside the bunker.
The large pancake shielding block has been split in two to allow for examination of residue found on the internal structure. Samples have been taken and analysis performed in the on-site chemistry lab.
After x-rays of the pipe welds to the shielding blocks, a decision has been made to lift some of the blocks out from the Monolith vessel to perform re-welds.
The base slab for the ESS Test Beamline has been cast. The concrete has cured and the formwork was removed.
ESTIA and SKADI
Base slabs were also cast for the two instruments occupying beamlines E1 and E2 – ESTIA and SKADI - in the east sector inside Experimental Hall 1.
Installation of casing and neutron optics continue in the E02 Beamline Gallery. Most of the casing is in place and neutron guides have been inserted in about 50% of the stretch from bunker to cave. Vacuum tests are ongoing for the section closest to the cave.
Neutron Beam Port Inserts
All eight inserts for the north, south and east sectors have been equipped with neutron optics and are sealed. The VESPA insert is last out to be vacuum tested, then all eight are ready for installation.
BIFROST, CSPEC, NMX and MAGIC inserts will follow in December.
North sector installations will be first off for both Neutron Beam Port Inserts and Plugs (NBPI and NBPP), but can only start once the Target inner shielding installation is completed to the matching level.
Neutron shielding bunker
Piping is under installation for the Neutron Beam Optics Assembly (NBOA) cooling, positioned within each Light Shutter System frame.